Find the first six terms of the sequence. a1 = -4, an = an-1 + 7


1
Expert's answer
2021-02-18T13:49:46-0500

"a_1=-4"


"a_2=a_1+7=-4+7=3"


"a_3=a_2+7=3+7=10"

"a_4=a_3+7=10+7=17"


"a_5=a_4+7=17+7=24"


"a_6=a_5+7=24+7=31"